OpenStudy (anonymous):

A=1 give your answers for A in the interval 0

OpenStudy (anonymous):

A=1 degrees

OpenStudy (anonymous):

for 2secA

OpenStudy (anonymous):

search on the net?

OpenStudy (anonymous):

It doesn't work.

OpenStudy (anonymous):

Is this the full question?

OpenStudy (anonymous):

No. 4=2secA (which is 1 degrees ~ because it's asking for degrees). Then you need to find the solutions in the intervals that I gave.

OpenStudy (anonymous):

ans is 60 degrees

OpenStudy (anonymous):

Why?

OpenStudy (anonymous):

4 = 2sec(A) divide both sides by 2 2 = sec(A) sec(A) = [1/cos(A)] For this reason multioly both sides by cos(A), OR u can say cross multiply 2cos(A) = 1 Divide both sides by 2 cos(A) = 1/2 A = \[\cos^{-1} (1/2)\] So A = 60

OpenStudy (anonymous):

Hm Ok. But wy first switch the 2 to 1/2 because adding the cos? Which would give 1...

OpenStudy (anonymous):

Which step are u talking about

OpenStudy (anonymous):

Um, never mind. I get it. :)

OpenStudy (anonymous):

There's another answer though. The other is 300.

OpenStudy (anonymous):

Cool, good luck

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Your welcome, how is your calculus coming along

OpenStudy (anonymous):

cos function repeats after 2(pi) degrees hence general solutions for cosine functions are \[2n \pi -\alpha\] where alpha is the angle hence put n=1 hence you get your other answer as 300 degrees